# Attribution Analysis Meets Model Editing: Advancing Knowledge Correction in Vision Language Models with VisEdit

### v1

@@ 1 Introduction
The Proposed VLLM Editor VisEdit: According to the above results, we design an effective VLLM editor. Based on the first attribution experiment, we place a trainable visual representation adaptor before the high-contribution layer. The adaptor applies cross-attention to infuse information of the edit sample into the visual representations of a given input sample. To ensure that the adaptation of visual representations is applied to the most important regions, based on the second experiment, we introduce an influence mapper module to identify the key visual regions most relevant to the edit prompt. In this way, the editor will focus on the visual representations that are crucial for modifying key responses, thereby enhancing the efficacy of editing while leaving irrelevant visual representations mostly untouched. We conduct editing experiments on three typical VLLMs, including BLIP2-OPT (2.7B) (Li et al. 2023), MiniGPT-4 (7B) (Zhu et al. 2023), and LLaVA-V1.5 (7B) (Liu et al. 2023), with two VLLM editing datasets, E-VQA and E-IC (Cheng et al. 2023). The experimental results demonstrate that our method excels in reliability, text/modal generality, and text/modal locality metrics, which follow standard evaluation protocols of the benchmark (Cheng et al. 2023). Furthermore, comprehensive exploratory experiments have validated the effectiveness of our module designs.
@@ 2.2 Attribution Analysis
Attribution analysis is an important tool to enhance the interpretability of deep learning models (Guidotti et al. 2019), where we primarily focus on perturbation-based works in this section. For example, (Ribeiro et al. 2016; Fong et al. 2017) align the predictions of the interpreter by randomly masking or perturbing image sections to determine the contributions of different image parts. By integrating feature perturbation with causal mediation analysis (Pearl 2001; Vig et al. 2020), Meng et al. (2022) identify the layers in LLMs that contribute most to the responses. For visual attribution analysis of VLLMs, (Schwettmann et al. 2023; Pan et al. 2023) use attribution methods such as integrated gradients (Sundararajan et al. 2017) to locate neurons in VLLMs that recognize specific visual concepts. Basu et al. (2024) demonstrate that the last few visual representations significantly influence the visual constraints of the prompt at the first model layer. However, these works do not consider how the spatial semantics of visual representations influence VLLM prediction.
@@ 3 Attribution Analysis
We evaluate the impact of the visual representations on token prediction in two steps. First, we compute the contribution of each layer’s module output to the predicted key token. Then, for a specific layer, we employ a noise perturbation method to evaluate how changes in a visual hidden state (synonymous with representation in the context) affect the attention module output. Figure 1 displays the data flow of VLLM response generation and the results of two attribution experiments. For more attribution analysis results, please refer to Appendix B.
Given a VLLM $f_{\theta}:\mathcal{X}_{v}\times\mathcal{X}_{t}\mapsto\mathcal{O}$ that maps an image-prompt pair $(x_{v},x_{t})$ into a text response $o=f_{\theta}(x_{v},x_{t})$, we set $\hat{f}_{\theta}:\mathcal{E}_{v}\times\mathcal{E}_{t}\mapsto\mathcal{Y}$ as the transformer in $f_{\theta}$ that maps an embedding sequence $\boldsymbol{\varepsilon}=\boldsymbol{\varepsilon}_{v}\oplus\boldsymbol{\varepsilon}_{t}\in\mathcal{R}^{N\times d_{h}}$ to a probability distribution $y\in\mathcal{Y}\subset\mathcal{R}^{|\mathcal{V}|}$ predicting the next token based on vocabulary $\mathcal{V}$, where $\boldsymbol{\varepsilon}_{v}\in\mathcal{E}_{v}\subset\mathcal{R}^{N_{v}\times d_{h}}$ is the image embedding and $\boldsymbol{\varepsilon}_{t}\in\mathcal{E}_{t}\subset\mathcal{R}^{N_{t}\times d_{h}}$ is the prompt embedding. $N_{v},N_{t},N$ are the length of visual, text, and complete embeddings, respectively. $d_{h}$ is the middle dimension of $\hat{f}_{\theta}$ and $\oplus$ denotes concatenation. In our attribution experiments, we represent each sample as a tuple $(\boldsymbol{\varepsilon}_{v},\boldsymbol{\varepsilon}_{t},o^{*})$, containing the image embedding $\boldsymbol{\varepsilon}_{v}$, the prompt embedding $\boldsymbol{\varepsilon}_{t}$, and the key token $o^{*}$ to be predicted.
@@ 3.2 Visual Representation Attribution
The heatmaps in Figure 1 display the visualization results (For more results, please refer to Appendix B.2). It can be observed that the model focuses on areas highly relevant to the prompt at deep layers.
@@ 4.1 Preliminaries
$\mathbb{E}_{(x^{e}_{v},x^{e}_{t},o^{e})\sim\mathcal{D}_{e}}\mathbb{I}\left\{f_{\theta_{e}}(x^{e}_{v},x^{e}_{t})=o^{e}\right\}$ (9)
@@
$\mathbb{E}_{(x^{e}_{v},x^{e}_{t},o^{e})\sim\mathcal{D}_{e}}\mathbb{E}_{x^{mg}_{v}\sim\mathcal{D}_{mg}(x^{e}_{v})}\mathbb{I}\left\{f_{\theta_{e}}(x^{mg}_{v},x^{e}_{t})=o^{e}\right\}$ (10)
$\mathbb{E}_{(x^{e}_{v},x^{e}_{t},o^{e})\sim\mathcal{D}_{e}}\mathbb{E}_{x^{tg}_{t}\sim\mathcal{D}_{tg}(x^{e}_{t})}\mathbb{I}\left\{f_{\theta_{e}}(x^{e}_{v},x^{tg}_{t})=o^{e}\right\}$ (11)
@@
$\mathbb{E}_{(x^{e}_{v},x^{e}_{t},o^{e})\sim\mathcal{D}_{e}}\mathbb{E}_{(x^{ml}_{v},x^{ml}_{t},o^{ml})\sim\mathcal{D}_{ml}(x^{e}_{v},x^{e}_{t})}$ (12)
@@
$\mathbb{E}_{(x^{e}_{v},x^{e}_{t},o^{e})\sim\mathcal{D}_{e}}\mathbb{E}_{(x^{tl}_{t},o^{tl})\sim\mathcal{D}_{tl}(x^{e}_{t})}$ (13)
@@ 5.1 Experimental Settings
Baseline Editors: To the best of our knowledge, there are currently no editors specifically designed for VLLMs. Therefore, most existing works (Cheng et al. 2023; Li et al. 2024) applied LLM editors in the VLLM case, including FT-V (Fine-tunes visual encoder), TF-L (Fine-tunes last layer of the language model), KE (Cao et al. 2021), IKE (Zheng et al. 2023), SERAC (Mitchell et al. 2022b), MEND (Mitchell et al. 2022a), TP (Huang et al. 2023), and LTE (Jiang et al. 2024). For experimental setup details, as well as the model hyperparameters and training specifics, please refer to Appendix A.
@@ 5.2 Analysis of Editing Performance
The overall editing performance is exhibited in Table 1 and Appendix C.1. Here we analyze Table 1 from different perspectives.
@@ 5.3 Analysis of VEAD Internals
Visualization of IM Module: We visualize the outputs of IM in Figure 3 (For more results and analysis, please refer to Appendix C.2). Specifically, Test 1 and Test 2 indicate that IM indeed picks the relevant visual region utilizing the overall semantics of the edit prompt, even when the visual objects in the edited samples are not entirely consistent with those in the input samples, e.g., the peeled and unpeeled bananas in test 2. Test 3 demonstrates that the visual areas of samples unrelated to the edit sample will receive low edit intensities from IM, because of $\ell_{im_{a}}$ and $\ell_{im_{\downarrow}}$.
Instance Analysis for VEAD: Figure 4 illustrates a counterfactual edit example where VEAD forces LLaVA-V1.5 to follow the knowledge even if it is incorrect (For more results and analysis, please refer to Appendix C.3). We first compare the visualization results of three cases: reliability samples before editing (BE Rel.), reliability samples after editing (AE Rel.), and generality samples after editing (AE T/M-Gen.). We observe that the editing process does not significantly change the visual regions VLLM focuses on for similar prompts, even if the “skateboarding” is counterfactually edited to “football”. The above observation indicates that VEAD adapts the visual representations in the skateboard region to new visual representations for football, and furthermore alters the final prediction to football, thereby validating the design objective of our model. Additionally, the visualization of modal locality samples shows that VEAD maintains VLLM’s original attention to visual objects unrelated to the edit.
@@ 6 Conclusion
This paper proposes a two-step attribution method based on contribution allocation and noise perturbation to measure the impact of visual representations on token prediction in VLLM. The attribution results show that the mid-to-late layers of VLLM pay particular attention to visual regions highly relevant to the prompt semantics. Based on the attribution results, we propose a novel VLLM editor, VisEdit. A visual editing adapter is inserted before the high-contribution layers to integrate editing information into the visual representations. Inside the editor, a novel IM module is introduced to enhance the edit signal in the most relevant visual regions and ignore the irrelevant ones. In VLLM editing experiments, VisEdit demonstrates exceptional editing performance. For future work, we plan to conduct further research on the attribution results of VLLM and explore expanding VEAD to multiple editing scenarios, such as by incorporating retrieval mechanisms.
Appendix A Details of Experimental Settings
A.1 Datasets:
@@ A.3 Baseline Editors:
A.4 Model Settings and Training Details:
@@
Appendix B Additional Attribution Analysis
B.1 Module Contribution Attribution
@@
B.2 Visual Representation Contribution Attribution
@@
Appendix C Additional Experimental Results
@@
C.1 Editing Performance on MiniGPT-4
@@ C.2 Results of IM Visualization
C.3 Instance Analysis for VEAD

### v3

@@ 1 Introduction
The Proposed VLLM Editor VisEdit: According to the above results, we design an effective VLLM editor. Based on the first attribution experiment, we place a trainable visual representation adaptor before the high-contribution layer. The adaptor applies cross-attention to infuse information of the edit sample into the visual representations of a given input sample. To ensure that the adaptation of visual representations is applied to the most important regions, based on the second experiment, we introduce an influence mapper module to identify the key visual regions most relevant to the edit prompt. In this way, the editor will focus on the visual representations that are crucial for modifying key responses, thereby enhancing the efficacy of editing while leaving irrelevant visual representations mostly untouched. We conduct editing experiments on three typical VLLMs, including BLIP2-OPT (2.7B) (Li et al. 2023), MiniGPT-4 (7B) (Zhu et al. 2023), and LLaVA-V1.5 (7B) (Liu et al. 2023), with two VLLM editing datasets, E-VQA and E-IC (Cheng et al. 2023). The experimental results demonstrate that our method excels in reliability, text/modal generality, and text/modal locality metrics, which follow standard evaluation protocols of the benchmark (Cheng et al. 2023). Furthermore, comprehensive exploratory experiments have validated the effectiveness of our module designs. Source codes are available at https://github.com/qizhou000/VisEdit.
@@ 2.2 Attribution Analysis
Attribution analysis enhances the interpretability of deep learning models (Guidotti et al. 2019), with this section focusing on perturbation-based methods. For instance, Ribeiro et al. (2016); Fong et al. (2017) use random masking or perturbation to assess contributions of image parts. Combining feature perturbation with causal mediation analysis (Pearl 2001; Vig et al. 2020), Meng et al. (2022) identify LLM layers that significantly influence responses. For visual attribution in VLLMs, methods like integrated gradients (Sundararajan et al. 2017) are used to locate neurons recognizing specific visual concepts (Schwettmann et al. 2023; Pan et al. 2023). Basu et al. (2024) demonstrate that the last few visual representations significantly influence the visual constraints of the prompt at the first model layer. However, these works do not consider how the spatial semantics of visual representations influence VLLM prediction.
@@ 3 Attribution Analysis
We evaluate the impact of the visual representations on token prediction in two steps. First, we compute the contribution of each layer’s module output to the predicted key token. Then, for a specific layer, we employ a noise perturbation method to evaluate how changes in a visual hidden state (synonymous with representation in the context) affect the attention module output. Figure 1 displays the data flow of VLLM response generation and the results of two attribution experiments. For more attribution analysis results, please refer to Appendix C.
Given a VLLM $f_{\theta}:\mathcal{X}_{v}\times\mathcal{X}_{t}\mapsto\mathcal{O}$ that maps an image-prompt pair $(x_{v},x_{t})$ into a text response $o=f_{\theta}(x_{v},x_{t})$, we set $\hat{f}_{\theta}:\mathcal{E}_{v}\times\mathcal{E}_{t}\mapsto\mathcal{Y}$ as the transformer in $f_{\theta}$ that maps an embedding sequence $\boldsymbol{\varepsilon}=\boldsymbol{\varepsilon}_{v}\oplus\boldsymbol{\varepsilon}_{t}\in\mathcal{R}^{N\times d_{h}}$ to a probability distribution $y\in\mathcal{Y}\subset\mathcal{R}^{|\mathcal{V}|}$ predicting the next token based on vocabulary $\mathcal{V}$, where $\boldsymbol{\varepsilon}_{v}\in\mathcal{E}_{v}\subset\mathcal{R}^{N_{v}\times d_{h}}$ is the image embedding and $\boldsymbol{\varepsilon}_{t}\in\mathcal{E}_{t}\subset\mathcal{R}^{N_{t}\times d_{h}}$ is the prompt embedding. $N_{v},N_{t},N$ are the length of visual, text, and complete embeddings, respectively. $d_{h}$ is the middle dimension of $\hat{f}_{\theta}$ and $\oplus$ denotes concatenation. In our attribution experiments, we represent each sample as a tuple $(\boldsymbol{\varepsilon}_{v},\boldsymbol{\varepsilon}_{t},o^{*})$, containing the image embedding $\boldsymbol{\varepsilon}_{v}$, the prompt embedding $\boldsymbol{\varepsilon}_{t}$, and the key token $o^{*}$ to be predicted. The symbols and notations used in this paper are provided in Appendix A.
@@ 3.2 Visual Representation Attribution
The heatmaps in Figure 1 display the visualization results (For more results, please refer to Appendix C.2). It can be observed that the model focuses on areas highly relevant to the prompt at deep layers.
@@ 4.1 Preliminaries
$\mathbb{E}_{(x^{e}_{v},x^{e}_{t},o^{e})\sim\mathcal{D}_{e}}\mathbb{I}\left\{f_{\theta_{e}}(x^{e}_{v},x^{e}_{t})=o^{e}\right\}$
@@
$\mathbb{E}_{(x^{e}_{v},x^{e}_{t},o^{e})\sim\mathcal{D}_{e}}\mathbb{E}_{x^{mg}_{v}\sim\mathcal{D}_{mg}(x^{e}_{v})}\mathbb{I}\left\{f_{\theta_{e}}(x^{mg}_{v},x^{e}_{t})=o^{e}\right\}$
$\mathbb{E}_{(x^{e}_{v},x^{e}_{t},o^{e})\sim\mathcal{D}_{e}}\mathbb{E}_{x^{tg}_{t}\sim\mathcal{D}_{tg}(x^{e}_{t})}\mathbb{I}\left\{f_{\theta_{e}}(x^{e}_{v},x^{tg}_{t})=o^{e}\right\}$
@@
$\mathbb{E}_{(x^{e}_{v},x^{e}_{t},o^{e})\sim\mathcal{D}_{e}}\mathbb{E}_{(x^{ml}_{v},x^{ml}_{t},o^{ml})\sim\mathcal{D}_{ml}(x^{e}_{v},x^{e}_{t})}$
@@
$\mathbb{E}_{(x^{e}_{v},x^{e}_{t},o^{e})\sim\mathcal{D}_{e}}\mathbb{E}_{(x^{tl}_{t},o^{tl})\sim\mathcal{D}_{tl}(x^{e}_{t})}$
@@ 5.1 Experimental Settings
Baseline Editors: To the best of our knowledge, there are currently no editors specifically designed for VLLMs. Therefore, most existing works (Cheng et al. 2023; Li et al. 2024) applied LLM editors in the VLLM case, including FT-V (Fine-tunes visual encoder), TF-L (Fine-tunes last layer of the language model), KE (Cao et al. 2021), IKE (Zheng et al. 2023), SERAC (Mitchell et al. 2022b), MEND (Mitchell et al. 2022a), TP (Huang et al. 2023), and LTE (Jiang et al. 2024). For experimental setup details, as well as the model hyperparameters and training specifics, please refer to Appendix B.
@@ 5.2 Analysis of Editing Performance
The overall editing performance is exhibited in Table 1 and Appendix D.1. Here we analyze Table 1 from different perspectives.
@@ 5.3 Analysis of VEAD Internals
Visualization of IM Module: We visualize the outputs of IM in Figure 3 (For more results and analysis, please refer to Appendix D.2). Specifically, Test 1 and Test 2 indicate that IM indeed picks the relevant visual region utilizing the overall semantics of the edit prompt, even when the visual objects in the edited samples are not entirely consistent with those in the input samples, e.g., the peeled and unpeeled bananas in test 2. Test 3 demonstrates that the visual areas of samples unrelated to the edit sample will receive low edit intensities from IM, because of $\ell_{im_{a}}$ and $\ell_{im_{\downarrow}}$.
Instance Analysis for VEAD: Figure 4 illustrates a counterfactual edit example where VEAD forces LLaVA-V1.5 to follow the knowledge even if it is incorrect (For more results and analysis, please refer to Appendix D.3). We first compare the visualization results of three cases: reliability samples before editing (BE Rel.), reliability samples after editing (AE Rel.), and generality samples after editing (AE T/M-Gen.). We observe that the editing process does not significantly change the visual regions VLLM focuses on for similar prompts, even if the “skateboarding” is counterfactually edited to “football”. The above observation indicates that VEAD adapts the visual representations in the skateboard region to new visual representations for football, and furthermore alters the final prediction to football, thereby validating the design objective of our model. Additionally, the visualization of modal locality samples shows that VEAD maintains VLLM’s original attention to visual objects unrelated to the edit.
@@ 6 Conclusion
This paper proposes a two-step attribution method combining contribution allocation and noise perturbation to measure the impact of visual representations on token prediction in VLLM. Results reveal that mid-to-late layers of VLLM focus on visual regions closely related to prompt semantics. Leveraging these insights, we introduce VisEdit, a novel VLLM editor with a visual editing adapter that integrates editing information into high-contribution layers. The editor incorporates an IM module to enhance relevant edit signals while ignoring irrelevant ones. VisEdit achieves outstanding performance in VLLM editing tasks. Future work will explore deeper insights into VLLM attributions and extend VEAD to broader editing scenarios.
Acknowledgments
This work is supported by the National Science and Technology Major Project (Grant No. 2022ZD0120302).
Appendix A Notations
The symbols and notations for variables and functions are presented in Table 4 and Table 5, respectively.
Variables Domain Explanation & Variants
$(x_{v},x_{t})$ $(\mathcal{X}_{v},\mathcal{X}_{t})$
an input image-text pair to a VLLM.
$x^{e}_{\cdot}$ is the input for editing.
$x^{mg}_{\cdot},x^{tg}_{\cdot}$ are the inputs for modal/text generality, respectively.
$x^{ml}_{\cdot},x^{tl}_{\cdot}$ are the inputs for modal/text locality, respectively.
$o$ $\mathcal{O}$
a text response from a VLLM.
$o^{*}$ is the key token to be predicted in attribution experiments.
$o^{e},o^{ml},o^{tl}$ are the predicting objects for editing and modal/text locality samples, respectively.
$\boldsymbol{\varepsilon}$ $\mathcal{R}^{N\times d_{h}}$
embedding of an input image-text pair.
$\boldsymbol{\varepsilon}_{v},\boldsymbol{\varepsilon}_{t}$ are the embeddings for vision and text input, respectively.
$y$ $\mathcal{R}^{\|\mathcal{V}\|}$ predicted probability distribution.
$N$ $\mathcal{Z}^{+}$
length of embeddings sequence.
$N^{\cdot}_{v},N^{\cdot}_{t},N^{\cdot}_{vt}$ are respectively the lengths of the vision, the text, and their concatenated hidden states.
$N^{e}_{\cdot}$ is the representation length of an editing sample.
$L$ $\mathcal{Z}^{+}$ the layer count of the transformer in the VLLM.
$l_{e}$ $[1,2,...,L]$ the index of layer to be edted.
$d_{h}$ $\mathcal{Z}^{+}$ hidden state dimension of the transformer in the VLLM
$h_{n}^{l}$ $\mathcal{R}^{N\times d_{h}}$
hidden state of the $n$-th token in the transformer at $l$-th layer.
$\tilde{h},\bar{h}$ are the perturbed hidden state and the hidden state of editing sample, respectively.
$\dot{h},\ddot{h}$ are the intermediate results in Influence Mapper.
$\acute{h},\grave{h},\breve{h}$ are the hidden states of reliability, generality, and locality samples, respectively.
$\mathring{h}$ is the hidden state randomly sampled from $\{\acute{h},\grave{h}\}$.
$m_{n}^{l}$ $\mathcal{R}^{N\times d_{h}}$ MLP output of the $n$-th token in the transformer at $l$-th layer.
$a_{n}^{l}$ $\mathcal{R}^{N\times d_{h}}$ Attention output of the $n$-th token in the transformer at $l$-th layer.
Table 4: Variable Symbols.
Functions Explanation & Variants
$f_{\theta}$
VLLM that maps input image-text pair to text
response, i.e., $o=f_{\theta}(x_{v},x_{t})$.
$\hat{f}_{\theta}$ is the language transformer in the VLLM.
$C_{o}(r)$ contribution of $r$ to $o$
$\delta$ softmax
$\sigma$ sigmoid
$\mathbb{E}$ expectation
ME model editor
$\oplus$ concatenation
Table 5: Function Symbols.
Appendix B Details of Experimental Settings
B.1 Datasets:
@@ B.3 Baseline Editors:
B.4 Model Settings and Training Details:
@@
Appendix C Additional Attribution Analysis
C.1 Module Contribution Attribution
@@
C.2 Visual Representation Contribution Attribution
@@
Appendix D Additional Experimental Results
@@
D.1 Editing Performance on MiniGPT-4
@@ D.2 Results of IM Visualization
D.3 Instance Analysis for VEAD